OpenStudy (gabrielah96):

write an equation of an ellipse with foci at (+-3,0) and vertices at (+-6,0)

OpenStudy (anonymous):

lol everyone hates these conic section problems they avoid them like the plague first off, is it clear what it looks like?

OpenStudy (anonymous):

we need that before we can start

OpenStudy (gabrielah96):

yes

OpenStudy (anonymous):

so you know that the larger number goes under the \(x^2\) term right?

OpenStudy (anonymous):

general form is \[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\] in your case it should be clear from the vertices an foci that the center is \((0,0)\) so it is just going to be \[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\]

OpenStudy (gabrielah96):

thats it?

OpenStudy (anonymous):

we need \(a\) and \(b\)

OpenStudy (anonymous):

you are pretty much given \(a\) in the question since the foci are \((-6,0)\) and \((6,0)\) you know \(a=6\) and so \(a^2=36\)

OpenStudy (anonymous):

now we are here \[\frac{x^2}{36}+\frac{y^2}{b^2}=1\] but we still need \(b\)

OpenStudy (gabrielah96):

and b?

OpenStudy (anonymous):

\[a^2-c^2=b^2\] and in this case since the foci are \((-3,0)\) and \((3,0)\) you get \(c=3\) making \[a^2-c^2=b^2\\ 36-9=b^2\\ 27=b^2\]

OpenStudy (anonymous):

final answer, assuming i did not make a mistake, is \[\frac{x^2}{36}+\frac{y^2}{27}=1\] want to check it ?

OpenStudy (gabrielah96):

how do i check it

OpenStudy (anonymous):

looks like we got the foci and vertices correct right? that is the easy check

OpenStudy (shiraz14):

@satellite73 : A typo again - in your previous post, you stated the following: "since the foci are (−6,0) and (6,0) you know a=6 and so a^2=36 " This should instead read as follows: "since the vertices are (−6,0) and (6,0) you know a=6 and so a^2 =36" I need to highlight these errors as someone new to this area (& wishes to learn it) might be confused with the comments you make. Thanks.